OpenStudy (anonymous):

Hello how would I factor 81x2 −9 y2 ? the x and y are squared

OpenStudy (anonymous):

what u need help on

OpenStudy (uri):

81x2 −9 y2 9xy(9x-y)

OpenStudy (anonymous):

9xy(9x-y)

OpenStudy (anonymous):

well\[81x^{2} -9y^{2} i need help with factoring 50 cent\]

OpenStudy (anonymous):

well how did you guys get that answer

OpenStudy (anonymous):

They factored it.

OpenStudy (uri):

Take the common.

OpenStudy (anonymous):

first they divided by nine

OpenStudy (uri):

9xy was the common

OpenStudy (anonymous):

for me to understand i need you to write it out for me please or take it step by step with me because i still do not understand

zepdrix (zepdrix):

9xy....? \[\large 9xy(9x-y) \qquad=\qquad 81x^2y-9xy^2\] That doesn't look like the equation he started with :\ hmm

zepdrix (zepdrix):

\[\large 81x^2-9y^2\]81 = 9*9, so let's rewrite it that way, \[\large 9\cdot9x^2-9y^2\] It appears they only thing these terms have in common is a 9. Divide a 9 out of each term,\[\large 9x^2-y^2\]Then we have to multiply each term by 9 to keep it balanced,\[\large 9(9x^2-y^2)\] Do you understand how factoring works Sunlight?

OpenStudy (anonymous):

Hello Zepdrix I factored both 81 AND 9

zepdrix (zepdrix):

Oh I made a boo boo... we have the `difference of squares`.. so we can simplify this down a bit further.

OpenStudy (anonymous):

and i got 9*9 xx - 3*3yy

zepdrix (zepdrix):

Ah yes very good :) So you can rewrite each term as a square.\[\large (9x)^2-(3y)^2\]

zepdrix (zepdrix):

And from here, we can apply a special rule.

zepdrix (zepdrix):

\[\large a^2-b^2=(a-b)(a+b)\]

OpenStudy (anonymous):

what is the a and the b suppose to stand for?

zepdrix (zepdrix):

I used different letters to show you a rule. We can apply that rule to our problem. So in our case, our a=9x, b=3y and we can break it down using the rule I posted above.

zepdrix (zepdrix):

\[\large \color{royalblue}{a}^2-\color{red}{b}^2=(a-b)(a+b)\]\[\large \color{royalblue}{(9x)}^2-\color{red}{(3y)}^2=?\]

OpenStudy (anonymous):

(9x-3y)(9x+3y)

OpenStudy (anonymous):

is that okay?

zepdrix (zepdrix):

Yay good job \c:/

zepdrix (zepdrix):

Subtracting squared terms, try to remember that rule. It comes up a lot as you advance in math classes.

OpenStudy (anonymous):

lol uhm okay but can i use the foiling method afterwards to work backwards?

zepdrix (zepdrix):

Yes, that's a great way to check your work.

OpenStudy (anonymous):

Hey I just did the foil and I got it yay thank you thank you :o))

zepdrix (zepdrix):

cool c:

OpenStudy (anonymous):

well okay good bye for now